OpenStudy (anonymous):

how to solve the inequality in terms of intervals (x+3)(x-8)(x+8)>and = 0

myininaya (myininaya):

(x+3)(x-8)(x+8)=0 when x=3 or when x=8 or when x=-8 ----|----|---|---- -8 3 8 you have four intervals to test

myininaya (myininaya):

oops x=-3 or x=8 or x=-8

myininaya (myininaya):

---|---|---|--- -8 -3 8 these are the intervals we will test: \[(-\infty,-8);(-8,-3);(-3,8);(8, \infty)\]

myininaya (myininaya):

\[f(x)=(x+3)(x-8)(x+8)\] \[f(-9)=(-9+3)(-9-8)(-9+8)=(-)(-)(-)=-\]\[f(-4)=(-4+3)(-4-8)(-4+8)=(-)(-)(+)=+\] \[f(0)=(0+3)(0-8)(0+8)=(+)(-)(+)=-\] \[f(9)=(9+3)(9-8)(9+8)=(+)(+)(+)=+\]

myininaya (myininaya):

we want f>=0 its 0 when w have -8,-3,8 and bigger than 0 when we have (-8,-3) or (8,inf) so the answer is \[[-8,-3] \cup [8,\infty)\]

OpenStudy (anonymous):

is question asking fx? In fact , I have not done math practice for 15 years.